OpenStudy (anonymous):

Help please what is the factored form of 2x^2+x-3 a) (2x+3)(x-1) b) (2x+1)(x-3) c) (2x-3)(x+1) d) (2x-1)(x+3)

OpenStudy (anonymous):

a

OpenStudy (anonymous):

Try the answers ! For example : d): (2x-1)(x+3)=2x^2+6x-x-3=2x^2+5x-3 k ?:)

OpenStudy (anonymous):

k :) got it

OpenStudy (anonymous):

k !

OpenStudy (anonymous):

\[2x^2+x-3\]First separate the first number into a product of two,\[2x\]\[x\]Then find two numbers that multiplied together give you the last number (-3),\[3* -1\]or\[-3*1\]Now, find which pair adds up to the coefficient in the middle (1), considering the 2 as a factor for one of them,\[3+(-1*2)=3-2=1\]So the pair that works is,\[3,-1\]Now add them to the two products you made in the first step,\[2x+3\]\[x-1\]Which is the answer for A.